OpenStudy (anonymous):

Customers at a phone company are given 2 plans to choose from. The first plan has a one time set up charge of $28 and charges .09 cents per minute. The second plan does not have a set up charge and costs .13 cents per minute. After how many minutes of calls will the 2 plans equal the same charge?

OpenStudy (ddcamp):

Cost of plan 1 , A(m) = 28 + 0.9m Cost of plan 2, B(m) = 0 + 0.13m For what m does A(m) = B(m) HINT: Set the two equations equal to each other.